OpenStudy (mathmath333):

Find the number positive integer solutions of 4x+6y=200

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{Find the number positive integer solutions of}\hspace{.33em}\\~\\ & 4x+6y=200 \end{align}}\)

OpenStudy (mathmath333):

I need to find quick way

OpenStudy (empty):

I don't really know but you should look into the Euclidean Algorithm I think that might be a good path to start looking into.

OpenStudy (anonymous):

as this is Diophantine equation i agree with @Empty

OpenStudy (anonymous):

so lets try 2x+3y=100

ganeshie8 (ganeshie8):

yeah both equations are same but we don't really need euclid algorithm here as finding a particular solution seems kinda easy by inspection. By inspection, (50, 0) is one particular solution

OpenStudy (anonymous):

but Euclid would help in finding general solution to show it's infinite solutions right ?

ganeshie8 (ganeshie8):

euclid is for finding particular solution since we already figured that (50, 0) is a particular soltion, we can avoid euclid

OpenStudy (anonymous):

:3

OpenStudy (mathmath333):

I found out (50,0),(47,2),.....(2,32) but that took more than 2 min

ganeshie8 (ganeshie8):

finding "one" particular solution is enough

OpenStudy (anonymous):

anyway i like this way to show its infinite:- 

ganeshie8 (ganeshie8):

after you have one particular solution, try finding the "null" solution : 2x + 3y = 0 Easy to see that (-3, 2) solves above equation. Therefore the complete solution is given by : `(50, 0) + t(-3, 2)`

ganeshie8 (ganeshie8):

since you want just the positive integer solutions, solve : 50 - 3t > 0 0 + 2t > 0

OpenStudy (mathmath333):

50/3=16.

OpenStudy (mathmath333):

answer is 17

ganeshie8 (ganeshie8):

right, solve it simultaneously you should get an interval of "t" as solution

OpenStudy (anonymous):

aha i haven't note positive :O

ganeshie8 (ganeshie8):

I think the answer should be 16

OpenStudy (mathmath333):

0<t<16

ganeshie8 (ganeshie8):

nope 0 < t < 16.66 there are exactly 16 positive integers in that interval

OpenStudy (mathmath333):

0<t<17

ganeshie8 (ganeshie8):

0 < t < 50/3 leave it like that

OpenStudy (mathmath333):

ok u r right (50,0) doesn't count answer is 16

ganeshie8 (ganeshie8):

Yep! lets do one more example maybe ?

ganeshie8 (ganeshie8):

Find the number of positive integer solutions to the equation 7x + 13y = 700

ganeshie8 (ganeshie8):

If you prefer, here are the steps : 1) Find any one `particular` solution by inspection 2) Find the `null` solution 3) Write out the complete solution : `particular` + `null`

OpenStudy (mathmath333):

(100,0)

ganeshie8 (ganeshie8):

Yep, keep going

OpenStudy (mathmath333):

how to find null soln

ganeshie8 (ganeshie8):

as the name says, it is the solution to the equation 7x + 13y = 0

ganeshie8 (ganeshie8):

give it a try.. it would feel awesome if you figure out a method to find the null solution on ur own..

OpenStudy (mathmath333):

(-13,-7)

ganeshie8 (ganeshie8):

Very close, but no. plug them in and see if they really produce 0

OpenStudy (mathmath333):

(-13,7)

ganeshie8 (ganeshie8):

Excellent! that is one null solution. Notice that any multiple of that also works, so all the null solutions are given by ` t(-13, 7)` where ` t ` belongs to the set of integers

OpenStudy (mathmath333):

100-13t>0 0+7t>0

ganeshie8 (ganeshie8):

Yes, you have skipped step3 but ok..

ganeshie8 (ganeshie8):

go ahead find the total count

OpenStudy (mathmath333):

what is step 3 ?

OpenStudy (mathmath333):

oh this one "3) Write out the complete solution : particular + null"

ganeshie8 (ganeshie8):

Yes, I was refering to that..

OpenStudy (mathmath333):

7 is answer

ganeshie8 (ganeshie8):

Yep! congratulations! Now you know how to solve any linear diophantine equation of form \(ax+by=c\)

OpenStudy (mathmath333):

cool!

ganeshie8 (ganeshie8):

so what was your method for finding null solution ?

OpenStudy (mathmath333):

ax + by = 0 (-b,a)

ganeshie8 (ganeshie8):

thats it!

ganeshie8 (ganeshie8):

that works always! so finding null solution is trivial as you can seethe only hard part is finding a particular solution

OpenStudy (anonymous):

6y=200-4x 3y=100-2x \[y=\frac{ 100-2x }{ 3 }\] by hit and trial when x=2 \[y=\frac{ 100-4 }{ 3 }=32\] add successively 3 to the value of x x=2+3=5,y=30 x=5+3=8,y=28 x=11,y=26 .... x=47,y=2 x=50,y=0

OpenStudy (anonymous):

.